OpenStudy (anonymous):

Expanding Form (5r-4t)^4

OpenStudy (anonymous):

You have to use here Binomial Theorem...

OpenStudy (anonymous):

i know i just cant seem to figure it out :/

OpenStudy (anonymous):

Do you know about this theorem??

OpenStudy (anonymous):

i think its (a-b)^2=a^2-2ab+b^2

OpenStudy (anonymous):

But here, you have to use it for exponent 4..

OpenStudy (anonymous):

\[ (x-y)^4 = x^4 -4 x^3y +6 x^2 y^2 - 4 x y^3 + y^4. \]

OpenStudy (anonymous):

Follow the steps posted above.

OpenStudy (anonymous):

Any problem you are facing??

OpenStudy (anonymous):

\[(x+y)^n = {^nC_0}(x^n y^0) + {^nC_1}(x^{n-1}y^1) + {^nC_2}(x^{n-2}y^2) + \cdots + {^nC_ n}(x^0 y^n)\]

OpenStudy (anonymous):

hi, yea i recieved the answer already, i appreciate your time and effort for helping me out.

OpenStudy (anonymous):

Here when you put n = 4, x = 5r and y = -4t: \[(5r - 4t)^4 = ^4C_0(5r)^4 + ^4C_1(5r)^{3}(-4t)^1 + + ^4C_2(5r)^2(-4t)^2 + continued \] + ^4C_3((5r)^1)(-4t)^3 + ^4C_4(-4t)^4

OpenStudy (anonymous):

See, there is nothing in the answer, just learn how we solve each and every problem, what is the method we are using, that will help you a lot, then you can solve any type of problem you are given with... So, just chase the solution and not answer...